MICROWAVE AND AB INITIO INVESTIGATION OF (CH_{3})_{3}N⋯HF⋯HF

\underline{C. S. BRAUER}, M. B. CRADDOCK, S. W. HUNT, and K. R. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455; JENS UWE-GRABOW, Lehrgebiet Physikalische Chemie A, Institut für Physikalische Chemie und Elektrochemie, Universtät Hannover, Callinstr. 3-3a, D-30167 Hannover.

Rotational spectroscopy and ab initio calculations have been used to examine the effect of a single HF solvent molecule on the gas phase proton transfer in the hydrogen bonded complex (CH_{3})_{3}N⋯HF. The rotational spectra of (CH_{3})_{3}N⋯HF⋯HF and five of its isotopically substituted derivatives have been observed by Fourier transform microwave spectroscopy. This follows a previous study by our group on H_{3}N⋯HF⋯HF, in which the addition of the second HF decreased the N⋯H hydrogen bond distance in H_{3}N⋯HF by 0.21(6) Å. The present study investigates the effect of the increased basicity of the amine on the N⋯H hydrogen bond. We observe a simple asymmetric rotor spectrum with strong a- and b- type transitions, consistent with a ring structure for the N⋯HF⋯HF frame. No evidence of internal rotation is observed. Structural analysis is underway and will be discussed, but preliminary analysis indicates an N⋯H hydrogen bond distance of about 1.4 Å, approximately halfway between the hydrogen bond distance in H_{3}N⋯HF of 1.7 Å and the N-H covalent length of 1.1 Å. Ab initio calculations concur with experiment, indicating that, as with H_{3}N⋯HF⋯HF, the trimethylamine complex forms a ring in which both the N⋯H hydrogen bond and the HFH angle are significantly perturbed. In addition, two of the methyl groups of the trimethylamine assume an eclipsed conformation, apparently participating in a bifurcated interaction with the fluorine atom of the second HF. This complex provides the first step in microsolvation of (CH_{3})_{3}N⋯HF and is useful in understanding the role of local environment in promoting proton transfer.